OpenStudy (anonymous):

how much work must be done to stop a 1500kg car traveling at 110km/hr?

OpenStudy (anonymous):

work must be equal to the kinetic energy of the car: \[E=mv^2/2\] you need to convert km/hr to m/s, (x1000/3600)

OpenStudy (anonymous):

Yah \[W = \Delta K = \cancel{K_f}-K_i=-K_i\]

OpenStudy (anonymous):

work done=energy. energy here is kinetic energy because of the presence of velocity. 1st convert 110km/hr to o.11m/s. 2rd E=1/2MV^2. SO 1/2*1500*(0.11)^2 7500*0.0121 =90.75J

OpenStudy (anonymous):

E (kenetic)=1/2 m v^2=w---->E=1/2 (1500 kg)*(1000m/3600s)*110km/h---->W=22.9kj